OpenStudy (anonymous):

If a line has a slope of 3/8 what would be the slope of a line perpendicular to it?

hartnn (hartnn):

remember this golden rule : *product of slopes of perpendicular lines = -1*

hartnn (hartnn):

if m1 = 3/8, and you want to find m2 , then do m1*m2 =3/8 and find m2

geerky42 (geerky42):

You mean \(m_1*m_2 =-1 \)?

hartnn (hartnn):

oops, typo

hartnn (hartnn):

i meant (3/8)*m2 = -1 find m2 .-.

geerky42 (geerky42):

Do you know how to find m2? @beccabear3

OpenStudy (anonymous):

No

geerky42 (geerky42):

ok so we have \(\dfrac{3}{8}m_2=-1\) to solve for m2, we need to isolate it how can we get 3/8 to the other side?

OpenStudy (anonymous):

multiply it out somehow?

OpenStudy (anonymous):

or subtract?

hartnn (hartnn):

lets take only '8' first, 8 is DIVIDED by m2 and we don't want it on the side of m2 so we do REVERSE operation, what is reverse of division ?? (addition or multiplication or subtraction?)

OpenStudy (anonymous):

Multiplication

hartnn (hartnn):

i mean m2 is DIVIDED by 8 and yes, multiplication is correct, so multiply 8 on both sides and tell me what u get ?

OpenStudy (anonymous):

Multiply 3/8 by 8 and -1 by 8?

hartnn (hartnn):

thats correct, do that

OpenStudy (anonymous):

3 and -8?

hartnn (hartnn):

so u get 3m2 =-8 right ??

hartnn (hartnn):

now 3 is MULTIPLIED by m2, and we don't want 3 on the same side of m2 so we do reverse operation of multiplication, which is division! so divide both sides by 3

OpenStudy (anonymous):

so -8/3?

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

welcome ^_^